# Slaves to the Law of Large Numbers: An Asymptotic Equipartition Property for Perplexity in Generative Language Models

### v1

@@ 3.2 Ensemble of Auxiliary Models
Let ${\mathbf{W}}$ denote the pseudo-random sequence that generated our string ${\mathbf{Y}}$ from model M. Thus we have ${\mathbf{Y}}=m({\mathbf{X}},{\mathbf{W}})$. With high probability, ${\mathbf{W}}$ is a typical sequence. Consider the ensemble model $M_{ensemble}$ for model M with pseudo-random sequences ${\mathbf{W}}_{1},~{}{\mathbf{W}}_{2}$ and let ${\mathbf{Y}}_{1}=m({\mathbf{X}},{\mathbf{W}}_{1})$. The classical AEP in Proposition 2 for model $M^{\prime}_{{\mathbf{Y}}_{1}}$ selected by the ensemble model requires that random strings ${\mathbf{Y}}^{\prime}_{N}$ generated by model $M^{\prime}_{{\mathbf{Y}}_{1}}$ must almost always satisfy
$\lim_{N\rightarrow\infty}|l_{M^{\prime}_{{\mathbf{Y}}_{1}}}({\mathbf{Y}}^{\prime}_{N})-h_{M^{\prime}_{{\mathbf{Y}}_{1}}}(N)|=0$ (18)
Since ${\mathbf{W}}$ is almost always a typical sequence, we can set ${\mathbf{W}}_{1}={\mathbf{W}}_{2}={\mathbf{W}}$, in which case the output of the ensemble model is exactly the same as the output ${\mathbf{Y}}$ of model M. Setting ${\mathbf{W}}_{1}\equiv{\mathbf{W}}$ we get ${\mathbf{Y}}_{1}\equiv{\mathbf{Y}}$ and
$h_{M^{\prime}_{{\mathbf{Y}}_{1}}}(N)=h_{M^{\prime}_{{\mathbf{Y}}}}(N)\equiv h_{M}({\mathbf{Y}}_{N}).$ (19)
Finally setting ${\mathbf{W}}_{2}\equiv{\mathbf{W}}$ gives ${\mathbf{Y}}^{\prime}\equiv{\mathbf{Y}}$ and
$l_{M^{\prime}_{{\mathbf{Y}}_{1}}}({\mathbf{Y}}^{\prime}_{N})=l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}_{N})\equiv l_{M}({\mathbf{Y}}_{N}).$ (20)
Inserting (19, 20) into (18) proves the result. ∎
Proposition 3 asserts that with high probability a sufficiently long text string will be typical for the model that it was generated from. Note the crucial role in the above proof of the assumption that the pseudo-random sequence ${\mathbf{W}}$ is typical. We can take this construct further to establish some non-asymptotic results as well. We expect from (16) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ will be within a few standard deviations of its mean e.g. $h_{M}({\mathbf{Y}}_{N})\pm 3\lambda_{M}({\mathbf{Y}}_{N})$ even when $N$ is not large enough for asymptotics to kick in. Finally, we will show that the set of typical strings is vanishingly small compared to all grammatically correct strings.
Proposition 4.
Let $\mathcal{T}_{M}^{N}\subset\mathcal{T}^{N}$ denote the set of substrings ${\mathbf{Y}}_{N}$ of typical strings ${\mathbf{Y}}$ for a model M. The size of the “typical set” $\mathcal{T}_{M}^{N}$ is an exponentially vanishing fraction $02^{-\gamma N},~{}\gamma>$ 0 of all grammatically correct sequences.
Proof.
If we use top-k sampling, the model M generates $k$ possible grammatically correct values for each token $Y_{n}$. The number $T_{gram}$ of grammatically correct strings is then lower-bounded as $T_{gram}>k^{N}\equiv 2^{N\log_{2}k}$. Under the (mild) assumption that for a long string ${\mathbf{Y}}_{N}$, at least some fraction of the token distributions $p_{n}(y)$ are non-uniform, we have $h_{M}({\mathbf{Y}}_{N})\leq h_{max}<\log_{2}k$. By definition, the probability of model M generating string ${\mathbf{Y}}_{N}$ is $2^{-N~{}l_{M}({\mathbf{Y}}_{N})}$, so we have for the probability $P_{typ}$ of a typical string ${\mathbf{Y}}_{N}$ generated by model M:
$P_{typ}$ $\doteq 2^{-N~{}l_{M}({\mathbf{Y}}_{N})}\rightarrow 2^{-N~{}h_{M}({\mathbf{Y}}_{N})}>2^{-N~{}h_{max}}$ (21)
The total number of typical strings cannot exceed $\left|\mathcal{T}_{M}^{N}\right|\leq 2^{N~{}h_{max}}<T_{gram}~{}2^{-N\left(\log_{2}k-h_{max}\right)}$. ∎
Proposition 4 suggests that the typical sets of two different models, each set being vanishingly small, likely have zero overlap.
4 Experiments with Open-Source Language Model
@@
We performed a series of experiments to verify the ideas described in Section 3. The basic idea is to use a local instance of the pre-trained GPT-2 model (Radford et al., 2019) to evaluate the empirical perplexity of generating a given string using the model’s conditional probability distributions and compare the result with the log-perplexity of the string again calculated using the model’s probability distributions. Software code for the experiments is available at Bell and Mudumbai (2024).
4.1 Perplexity for GPT-2 of a String Generated by GPT-2
@@
Our first set of experiments consisted of repeatedly generating a long text string from GPT-2 and then evaluating its log-perplexity, empirical entropy and log-deviations for the same GPT-2 model.
@@
We initialized the prompt string and the seed of random number generator with fixed values to allow for later reproduction. We then used the top-k sampling with some fixed $k$ to generate the probability distribution for the next token. We normalized the probabilities of the top-k tokens to sum to unity. We then created and saved a table consisting of token ID, token string and selection probability of each of the top $k$ tokens. Finally, a random next token is chosen by sampling from the normalized distribution, and the new token is appended to the prompt to repeat the process for $N_{max}=512$ total tokens which is the maximum number allowed by the GPT-2 model.
@@
The random token $Y_{n}$ chosen at each step along with the probability distribution $p_{n}(y)$ from which they were chosen were saved into a json file for later processing. We ran this experiment many times for different seed values for the random number generator, different choices of $k$ for the top-k sampling and a number of different initial prompts.
@@
For each generated string, we can then analyze the information archived in the json file to test our claim in Proposition 3. Specifically, for each token $N\in 1\dots N_{max}$ we calculated the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the substring ${\mathbf{Y}}_{N}$ consisting of the first $N$ tokens of the generated string.
4.2 Perplexity for GPT-2 of an Arbitrary String
@@
We also used a slightly modified version of the procedure described in Section 4.1 to calculate the log-perplexity and empirical entropy of an arbitrary string (that was not generated by GPT-2) on the GPT-2 model. We first choose a text string that we wish to analyze. We fix an initial fragment of the string as our prompt ${\mathbf{X}}$. We tokenize the remaining portion of the string excluding the prompt using the GPT-2 tokenizer. Let $Y_{n},~{}n=1\dots N_{max}$ denote the resulting sequence of tokens.
@@
Stating with $n=1$, we list the probability distribution $p_{n}(y)$ of the next token $Y_{n}$ for GPT-2 using top-k sampling for the $n$’th token. We can then calculate the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ as in Section 4.1. However, to calculate the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, we also need the probability that GPT-2 will output the actual next token $Y_{n}$ from the string ${\mathbf{Y}}$ which may not be included in the list of top-$k$ tokens [^3] Note that our choice of top-k sampling is a compromise. The number of tokens in GPT-2 is quite large ($|\mathcal{T}|\approx 40,000$). If we exhaustively list the full probability distribution of all tokens, our experiment becomes very computationally expensive and slow.. When this happens, we replaced the lowest probability entry in the list of top-k tokens with the actual next token from our string along with its probability.

### v3

@@ 3 An Equipartition Property for Perplexity
We want to argue that $\lambda_{M}({\mathbf{Y}}_{N})$ is the standard deviation of $l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})$, and therefore asymptotically vanishing $\lambda_{M}({\mathbf{Y}}_{N})$ implies $l_{M}({\mathbf{Y}}_{N})\rightarrow h_{M}({\mathbf{Y}}_{N})$. This, however, is very wrong - for one thing $\lambda_{M}({\mathbf{Y}}_{N})$ is itself a random variable. In the sequel, we propose an Indifference Principle that allows us to carry through reasoning like the above to prove Proposition 3.1 and potentially other results. But first we present a direct proof of Proposition 3.1.
3.1 Formal Proof: an LLN for LLMs
We will use a version of the Weak Law of Large Numbers that applies to sequences of random variables with no restrictions on their dependence. Such a Law is presented in Appendix B (Proposition B.1). In fact, it turns out that a weaker version of Proposition B.1 namely Lemma B.2 is sufficient for our purposes.
Proof of Proposition 3.1..
Consider the sequence of conditional likelihood random variables $\{l_{n}\}$ defined as:
$l_{n}$ $\doteq-\log_{2}\Pr\left(Y_{n}|{\mathbf{Y}}_{n-1}\right)\equiv-\log_{2}\left(p_{n}(Y_{n})\right)$ (16)
We note that the random variables $Z_{n}\doteq l_{n}$ satisfy all of the conditions for Lemma B.2, if we impose the restriction that they are strictly bounded i.e. $|l_{n}|<L$, for some some finite $L$. This is equivalent to requiring that all non-zero token probabilities $p_{n}(y)$ are above some lower bound e.g. $p_{n}(y)\geq 10^{-10}$. In practice, this is easily satisfied by any practical LLM. Then from Lemma B.2, we have for any $\epsilon>0$:
$\lim_{N\rightarrow\infty}\Pr\left(\left|\frac{1}{N}\sum_{n=1}^{N}\left(l_{n}-H(p_{n})\right)\right|>\epsilon\right)$ $=0$ (17)
From the definition (16), we note that $\frac{1}{N}\sum_{n=1}^{N}l_{n}\equiv l_{M}({\mathbf{Y}}_{N})$, and likewise $\frac{1}{N}\sum_{n=1}^{N}H(p_{n})\equiv h_{M}({\mathbf{Y}}_{N})$. Combining these observations with (17) gives (15). ∎
3.2 Informal Proof: Roads not Traveled
We now present an alternative proof of Proposition 3.1 using a more informal argument. Specifically, we introduce the conceptual device of an auxiliary language model that will allow us to apply the traditional Law of Large Numbers to a sequence of dependent random variables.
Figure 1: Probability tree for the first 3 tokens of a language model M that selects each token $Y_{n}$ from a dictionary $\mathcal{T}$ of three tokens. Also shown in red is the probability tree of an auxiliary model $M^{\prime}_{\mathbf{S}}$ for a string ${\mathbf{S}}$ with $S_{1}=3,~{}S_{2}=1$ and so on.
3.2.1 Concept of an Auxiliary Language Model
Let us consider a fixed string ${\mathbf{S}}$. For this particular string ${\mathbf{S}}$ and model M, we have a fixed set of distributions $q_{n}(y)\doteq\Pr(S_{n}=y|{\mathbf{S}}_{n-1})$. Now define an auxiliary generative model $M^{\prime}_{\mathbf{S}}$ that generates a random string ${\mathbf{Y}}^{\prime}\doteq[Y_{1}^{\prime},~{}Y_{2}^{\prime},~{}\dots,~{}Y_{N}^{\prime},~{}\dots]$ where the tokens $Y_{n}^{\prime}$ are generated independently from the distributions $q_{n}(y)$. If we represent a language model M as a probability tree, an auxiliary model $M^{\prime}_{\mathbf{S}}$ represents one specific subtree corresponding to the string ${\mathbf{S}}$ that defines the model. This is illustrated in Fig. 1. The auxiliary model $M^{\prime}_{\mathbf{S}}$ is much simpler than the general language model M, because its output tokens $Y^{\prime}_{n}$ are generated independently of other tokens. Our informal proof of Proposition 3.1 is based on the following idea.
Proposition 3.2.
The Indifference Principle. For any random event $E$ defined entirely on the subtree of a particular string ${\mathbf{S}}$, $\Pr(E)$ will be the same for generative model M as for the auxiliary model $M^{\prime}_{\mathbf{S}}$.
The idea behind Proposition 3.2 is that while calculating $\Pr(E)$ for some event $E$, we can change any part of the probability tree that do not affect the event $E$ (“the roads not traveled”) to make our calculation easier e.g. by allowing us to work with an auxiliary language model $M^{\prime}_{\mathbf{S}}$ which generates strings of independent tokens. For our purposes, Proposition 3.2 applies to all events in the $\sigma$-algebra generated by the conditional likelihood random variables $l_{n}$ as defined in (16).
Informal Proof of Proposition 3.1..
Using the Proposition 3.2, we can evaluate the distribution of random variables $l_{n}$ assuming the tokens $Y_{n}$ are generated by the auxiliary model $M^{\prime}_{\mathbf{S}}$ instead of the original model M. Then the random variables $l_{m},~{}l_{n}$ are independent of each other and $\operatorname*{{E}}\left[l_{n}\right]\equiv H\left(q_{n}\right)$. The weak Law of Large Numbers for the sequence of independent zero-mean random variables $l_{n}-H\left(q_{n}\right)$ immediately gives us (17). ∎
Proposition 3.2 allows us to use traditional statistical laws that apply to sequences of independent random variables directly without having to prove bespoke versions of those laws such as Proposition B.1. For instance, we may be able to use Proposition 3.2 to formulate a version of the Central Limit Theorem for $\frac{1}{\lambda_{M}({\mathbf{S}}_{N})}\left(l_{M}({\mathbf{S}}_{N})-h_{M}({\mathbf{S}}_{N})\right)$. We delegate a full development of these ideas to future work and limit ourselves to one simple non-asymptotic result to illustrate the method.
3.3 Chebyshev Bound on Deviations
We note that for the auxiliary model $M^{\prime}_{\mathbf{S}}$, the variance of the conditional log likelihood random variable $l_{n}$ is $\lambda^{2}(q_{n})$ as defined in (13). Assuming independence of $l_{n}$ using Proposition 3.2, we can show that the variance of $l_{M}({\mathbf{S}}_{N})$ is $\lambda^{2}_{M}({\mathbf{S}}_{N})$ as defined in (14). We can then write using the Chebyshev’s Inequality (Cohen, 2015) for any $\alpha>0$:
$\Pr\left(\left|l_{M}({\mathbf{S}}_{N})-h_{M}({\mathbf{S}}_{N})\right|>\alpha\lambda_{M}({\mathbf{S}}_{N})\right)$ $\leq\frac{1}{\alpha^{2}}$ (18)
3.4 Typical Sets are Vanishingly Small
Let us now formally define the typical set $\mathcal{P}_{M}^{n}(\epsilon)\subset\mathcal{T}^{n}$ of typical strings ${\mathbf{Y}}_{n}$ for a model M with $\epsilon>0$:
$\mathcal{P}_{M}^{n}(\epsilon)$$\doteq\left\{{\mathbf{Y}}_{n}:\left|l_{M}({\mathbf{Y}}_{n})-h_{M}({\mathbf{Y}}_{n})\right|<\epsilon\right\}$ (19)
$\equiv\left\{{\mathbf{Y}}_{n}:\left|h_{M}({\mathbf{Y}}_{n})+\frac{1}{n}\log_{2}\left(\Pr({\mathbf{Y}}_{n})\right)\right|<\epsilon\right\}$
Proposition 3.1 asserts that with high probability a stochastic model M is constrained to only output typical strings from this set $\mathcal{P}_{M}^{n}(\epsilon)$ for any $\epsilon>0$ if we consider sufficiently long strings. This observation has profound consequences. We will now show one such consequence, namely that the typical set $\mathcal{P}_{M}^{n}(\epsilon)$ is very small:
Proposition 3.3.
The size of the typical set $\mathcal{P}_{M}^{n}$ is an exponentially vanishing fraction $02^{-\gamma n},~{}\gamma>$ 0 of all grammatically correct sequences.
The argument below closely tracks standard information-theoretic results (Cover & Thomas, 2005a) about the size of the typical set. We start by restating Proposition B.1 informally as:
$\Pr({\mathbf{Y}}_{n})$ $\equiv 2^{-nl_{M}({\mathbf{Y}}_{n})}\rightarrow 2^{-nh_{M}({\mathbf{Y}}_{n})}$ (20)
Let $T_{g}(n)$ be the total number of grammatically correct strings ${\mathbf{Y}}_{n}$ of length $n$. The average entropy $h_{M}({\mathbf{Y}})$ is a measure of the richness of the model that generated the string. Language models typically have a “temperature” parameter $T$ in the range $T\in[0,1]$ that can be used to tune the richness of the model’s output. Higher temperature means less “predictable” (i.e. less deterministic) and more “creative” (i.e. more random) outputs. This corresponds to higher values for $h_{M}({\mathbf{Y}})$.
Consider the maximally rich model of length $n$ that has the highest possible entropy $h_{M}({\mathbf{Y}}_{n})\equiv h_{max}(n)$ which corresponds to choosing all possible (grammatically correct) strings of length $n$ with equal probability [^3] Note that this does not mean that the token distributions $p_{m}(y)$ are uniform over all valid tokens at each step $m$; indeed tokens corresponding to richer sub-trees as shown in Fig. 1 should be chosen with higher probability to maximize the overall richness of the model’s output.. By definition, for the maximally rich model, the probability of choosing any valid string is $2^{-nh_{max}(n)}$ and this means the number of possible strings of length $n$ is:
$T_{g}(n)$ $\equiv 2^{nh_{max}(n)}$ (21)
We can take (21) as the defining relation for $h_{max}(n)$. In practical LLMs, even at the highest temperature setting i.e. $T=1$, the token probability is typically concentrated in a small number of tokens and the token probabilities are determined by far more complex rules than the richness of their sub-trees. Let us formalize this as an assumption.
Assumption 3.1.
Practical LLMs are not entropy-maximizing. Any language model M has an entropy gap $\Delta h>0$ such that the following holds for all sufficiently long strings:
$h_{M}({\mathbf{Y}}_{n})$ $<h_{max}(n)-\Delta h,~{}\forall{\mathbf{Y}}_{n}\in\mathcal{T}^{n}$ (22)
Proof of Proposition 3.3.
From (20), for $n\gg 1$, we have $\Pr({\mathbf{Y}}_{n})\equiv 2^{-nl_{M}({\mathbf{Y}}_{n})}\to 2^{-nh_{M}({\mathbf{Y}}_{n})}>2^{-n(h_{max}(n)-\Delta h)}$ where we used Assumption 3.1 in the last step. Since this holds for any typical string of length $n$, we must have for small $\epsilon>0$ and correspondingly large $n$:
$\left|\mathcal{P}_{M}^{n}(\epsilon)\right|$ $<2^{n\left(h_{max}(n)-\Delta h)\right)}\equiv 2^{-n\Delta h}\left|T_{g}(n)\right|$ (23)
Thus, the typical set $\mathcal{P}_{M}^{n}(\epsilon)$ represents an exponentially vanishing fraction of all grammatically correct strings. ∎
We emphasize that all we needed to establish Proposition 3.3 was Assumption 3.1 which only requires that the model assigns higher probability to some sentences than to others.
3.5 Over-Typical Sets are Vanishingly Smaller Still
The obverse of Proposition 3.3 is that almost all grammatically correct strings are non-typical for a given language model M. A string can be non-typical in two different ways depending on whether its perplexity is abnormally high (“under-typical”) or abnormally low (“over-typical”). Formally, we define the set $\mathcal{V}_{M}^{n}(\epsilon)$ of over-typical strings of length $n$ for a model M as:
$\mathcal{V}_{M}^{n}(\epsilon)$$\doteq\left\{{\mathbf{Y}}_{n}:\left(h_{M}({\mathbf{Y}}_{n})-l_{M}({\mathbf{Y}}_{n})\right)>\epsilon\right\}$
$\equiv\left\{{\mathbf{Y}}_{n}:\Pr({\mathbf{Y}}_{n})>2^{-n\left(h_{M}({\mathbf{Y}}_{n})-\epsilon\right)}\right\}$ (24)
Proposition 3.4.
The size of the over-typical set $\mathcal{V}_{M}^{n}(\epsilon)$ is an exponentially vanishing fraction of all grammatically correct sequences.
@@ 3.5 Over-Typical Sets are Vanishingly Smaller Still
For any ${\mathbf{Y}}_{n}\in\mathcal{V}_{M}^{n}(\epsilon)$, we have from (24) and (22):
$\Pr({\mathbf{Y}}_{n})$ $>2^{-nh_{max}(n)}\times 2^{n(\Delta h+\epsilon)}$ (25)
Therefore, we must have
$\left|\mathcal{V}_{M}^{n}(\epsilon)\right|$ $<\left|T_{g}(n)\right|2^{-n(\Delta h+\epsilon)}$
where we used the definition of $h_{max}$ in (21). ∎
We have shown that almost all long grammatically correct strings are under-typical for any language model M. We illustrate these ideas for a simple Markov language model in Appendix A.
4 Experiments with Open-Source Models
@@
We performed a series of experiments to verify the ideas described in Section 3. The basic idea is to use a local instance of a pre-trained open-source model to evaluate the empirical entropy of a given string using the model’s conditional probability distributions and compare the result with the log-perplexity of the string again calculated using the model’s probability distributions. We ran experiments on the smaller GPT-2 model (Radford et al., 2019) as well as the more recent, larger and more capable Llama 3.1 8B (Touvron et al., 2023) model.
4.1 Perplexity of Self-Generated Strings
@@
Our first set of experiments consisted of repeatedly generating a long text string from a model (GPT-2 or Llama 3.1 8B) and then evaluating its log-perplexity, empirical entropy and log-deviations for the same model.
@@
We initialized the prompt string and the seed of random number generator with fixed values to allow for later reproduction. We then used top-k sampling with $k=100$ to generate the probability distribution for the next token. We normalized the probabilities of the top-k tokens to sum to unity. We then created and saved a table consisting of token ID, token string and selection probability of each of the top $k$ tokens. Finally, a random next token is chosen by sampling from the normalized distribution, and the new token is appended to the prompt to repeat the process for $N_{max}$ total tokens. We want the number of tokens $N_{max}$ to be large enough to observe the asymptotic behavior of the perplexity of the generated string. In practice, $N_{max}$ is constrained by certain limitations of the models as we discuss in the sequel.
@@
The random token $Y_{n}$ chosen at each step along with the probability distribution $p_{n}(y)$ from which they were chosen were saved into a json file for later processing. We ran this experiment many times for different seed values for the random number generator, different choices of sampling distributions and a number of different initial prompts.
@@
For each generated string, we can then analyze the information archived in the json file to test our claim in Proposition 3.1. Specifically, for each token $N\in 1\dots N_{max}$ we calculated the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the substring ${\mathbf{Y}}_{N}$ consisting of the first $N$ tokens of the generated string.
4.2 Perplexity of Externally Generated Strings
@@
We also used a slightly modified version of the procedure described in Section 4.1 to calculate the log-perplexity and empirical entropy of an arbitrary string (that was not generated by the model itself) on the GPT-2 and Llama 3.1 8B models. We first choose a text string that we wish to analyze. We fix an initial fragment of the string as our prompt ${\mathbf{X}}$. We tokenize the remaining portion of the string excluding the prompt using the model’s tokenizer. Let $Y_{n},~{}n=1\dots N_{max}$ denote the resulting sequence of tokens.
@@
Stating with $n=1$, we list the probability distribution $p_{n}(y)$ of the next token $Y_{n}$ for the model using top-k with $k=100$ sampling for the $n$’th token. We can then calculate the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ as in Section 4.1. However, to calculate the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, we also need the probability that our model will output the actual next token $Y_{n}$ from the string ${\mathbf{Y}}$ which may not be included in the list of top-$k$ tokens [^4] Note that our choice of top-k sampling is a compromise. The number of tokens in GPT-2 or Llama is quite large ($|\mathcal{T}|\approx 40,000$). If we exhaustively list the full probability distribution of all tokens, our experiment becomes very computationally expensive and slow.. When this happens, we replaced the lowest probability entry in the list of top-k tokens with the actual next token from our string along with its probability.